OpenStudy (anonymous):

what is the approximate slope of the curve with equation f(x)=1 /x-1 @ x=1.2

myininaya (myininaya):

Is that f(x)=1/x-1 or is it suppose to be f(x)=1/(x-1)?

OpenStudy (sidsiddhartha):

slope means derivative so differentiate the function with respect to x first

OpenStudy (anonymous):

fx=1/x-1 at x=1.2

OpenStudy (sidsiddhartha):

yeah do u have any knowledge of derivatives?

myininaya (myininaya):

you could easily use the formal definition of derivative which is \[\frac{df}{dx}=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\] so if you are given \[f(x)=\frac{1}{x}-1 \text{ then what is } f(x+h)=?\] I'm asking this because our formula calls for it